# SuperM2M: Supervised and Mixture-to-Mixture Co-Learning for Speech Enhancement and Robust ASR

### v2

@@ 7 Evaluation Results
Table 6.7.2 and 7 respectively report $1$- and $6$-channel enhancement performance on the fifth microphone of the CHiME-4 simulated test data, and robust ASR performance on the official CHiME-4 test utterances (by using the Whisper Large v2 model for recognition). The two tables consist of exactly the same set of experiments, and differ only in the number of input microphones to the enhancement models.
On the simulated test data, purely supervised models (in row $1$a-$1$b) trained on the official simulated training data (denoted as S in the “Training data” column) produce large improvement over unprocessed mixtures (e.g., in row 1a of Table 6.7.2, $17.1$ vs. $7.5$ dB SISDR, and in 1a of Table 7, $22.2$ vs. $7.5$ dB SISDR), and achieve enhancement performance better than existing supervised models such as iNeuBe (based on TCN-DenseUNet) [13, 81], SpatialNet [23], USES [24] and USES2 [25] in both $1$- and $6$-channel cases. However, the ASR performance is much worse than directly using unprocessed mixtures for ASR, especially in multi-channel cases (e.g., in row $1$a and 00 of Table 6.7.2, $10.20$% vs. $7.69$% WER on REAL test data, and in row $1$a and 00 of Table 7, $53.23$% vs. $7.69$% WER on REAL test data). This degradation has been widely observed by existing studies [33, 43], largely due to the mismatches between simulated training and real-recorded test conditions, and the speech distortion incurred by enhancement.
In row $7$ of Table 6.7.2 and 7, we respectively show the results of our best performing SuperM2M models (without employing the speaker reinforcement technique described in Section 6.7.3) for $1$- and $6$-channel cases. By training on the official simulated and real data combined (denoted as S+R), SuperM2M obtains clearly better ASR results on the real test data than the purely-supervised models (in row $1$a-$1$b) and unprocessed mixtures (in row 00), and the enhancement performance on the simulated test data remains strong. These results indicate that SuperM2M can effectively learn from real data, has better generalizability to real data, and can perform enhancement with low distortion to target speech and high reduction to non-target signals.
@@ 7.2 Effects of Including Loss on Close-Talk Mixture in $\mathcal{L}_{\text{MC}}$
First, we report the results of not including close-talk microphone in the loss function. That is, in (5) we do not include the third term, $\mathcal{L}_{\text{MC},0}$, when training SuperM2M. We observe clear improvement in ASR performance on the real test data over the purely-supervised model in row $1$a (i.e., $7.02$% vs. $10.20$% in Table 6.7.2 and $4.46$% vs. $53.23$% in Table 7), indicating that close-talk mixtures are not must-have for our system.
Next, we include close-talk mixtures for training. In row $4$a-$4$g, the number of future taps for the close-talk microphone, $\ddot{J}$, is tuned to a value shared by all the training utterances. We enumerate $\ddot{J}$ from $1$ all the way up to $8$. In row $5$, we instead estimate $\ddot{J}$, on the fly, for each training mixture during training, using the technique described in Section 4.4. Compared with row $4$a-$4$g, row $5$ obtains better ASR performance on the real test data, indicating the effectiveness of the proposed way of estimating $\ddot{J}$.
In both Table 6.7.2 and 7, comparing row $5$ and $3$, we observe that including close-talk mixtures for training SuperM2M produces better ASR performance on the real data of CHiME-4 (e.g., $6.51$% vs. $7.02$% in Table 6.7.2 and $3.97$% vs. $4.46$% in Table 7).
@@
In row $6$a-$6$d of Table 6.7.2 and 7, we respectively report the results of UNSSOR [35] and M2M [40]. UNSSOR is trained in an unsupervised way directly on far-field mixtures by using the $\mathcal{L}_{\text{MC}}$ loss in (5) but without including the third term $\mathcal{L}_{\text{MC},0}$ (i.e., the loss on close-talk mixture). M2M is the same as SuperM2M but without the supervised learning branch. We can train UNSSOR and M2M on the real data alone (denoted as R), or on the simulated and real data combined (i.e., S+R) considering that SuperM2M is trained on S+R. When M2M is trained on S+R, the loss function on the simulated data does not include the loss on close-talk mixture, since CHiME-4 does not provide simulated close-talk mixtures.
@@
Comparing row $6$a-$6$d and $5$, we can see that SuperM2M obtains much better performance than UNSSOR and M2M, suggesting the effectiveness of the proposed co-learning mechanism. The low performance of M2M and UNSSOR is possibly because the noise in CHiME-4 typically consists of an unknown number of diffuse and directional sources. In this case, unsupervised algorithms with two hypothesized directional sources tend to get confused about which source should be the target speech source. This problem is naturally avoided by the supervised learning mechanism in SuperM2M, which models noise sources as a single, combined source.
@@ 7.4 Miscellaneous Results
In row $7$ of Table 6.7.2 and 7, we switch TF-GridNet from v1 to v2, perform iSTFT-STFT projection, and apply SNR augmentation. From row $5$ to $7$, although the ASR performance on the real test data gets better in Table 6.7.2 but worse in Table 7, in both tables better ASR performance is observed on the real validation set. We therefore use the system in row $7$ as our default system for the rest experiments of this paper.
@@
Row Systems Frontend SIMU REAL SIMU REAL
@@
2b 7 of Table 6.7.2 SuperM2M TFGridNetv2 1 $10$ 2.400 442 4772.4004424772.400\,442\,4772.400 442 477 1.640 915 9621.6409159621.640\,915\,9621.640 915 962 4.538 662 6824.5386626824.538\,662\,6824.538 662 682 2.400 859 4512.4008594512.400\,859\,4512.400 859 451
@@
4b 7 of Table 7 SuperM2M TFGridNetv2 2 $10$ 1.283 185 841.283185841.283\,185\,841.283 185 84 1.331 170 0281.3311700281.331\,170\,0281.331 170 028 1.881 770 6381.8817706381.881\,770\,6381.881 770 638 1.835 676 5841.8356765841.835\,676\,5841.835 676 584
@@
6b 7 of Table 7 SuperM2M TFGridNetv2 6 $10$ 0.833 333 3330.8333333330.833\,333\,3330.833 333 333 1.231 608 8351.2316088351.231\,608\,8351.231 608 835 1.372 805 3791.3728053791.372\,805\,3791.372 805 379 1.578 775 2811.5787752811.578\,775\,2811.578 775 281
@@
Notes: The best scores are highlighted in bold in the $1$- and $6$-channel cases separately.
@@ 7.6 Breaking Out to New Highs on CHiME-4 ASR Tasks
Table 7.5 reports the ASR performance of SuperM2M based on the official ASR setup of CHiME-4, using the ASR model proposed in [53] and still following the evaluation pipeline in Fig. 3. Comparing row 00b with 00a, we observe that we have successfully reproduced the ASR system proposed in [53].
SuperM2M, despite not jointly trained with the ASR model, achieves a new state-of-the-art on the REAL test set in each of the $1$-, $2$- and $6$-channel tasks, significantly outperforming the previous best obtained by IRIS [53] and MultiIRIS [84] (e.g., in the $1$-channel case $3.04$% vs. $3.92$% WER in $2$a and $1$b, in the $2$-channel case $1.94$% vs. $2.65$% WER in $4$a and $3$b, and in the $6$-channel case $1.61$% vs. $1.77$% WER in $6$a and $5$b). IRIS jointly trains a Conv-TasNet based monaural enhancement model, a WavLM based ASR feature extractor, and an encoder-decoder transformer based ASR model. MultiIRIS, building upon IRIS, replaces the Conv-TasNet module with a DNN based weighted power minimization distortionless response (WPD) beamformer. From row $1$a vs. $1$b, $3$a vs. $3$b, and $5$a vs. $5$b, we observe that, without joint training, IRIS and MultiIRIS often obtain clearly worse ASR performance, especially in the $1$- and $2$-channel cases. These results further indicate the effectiveness and potential of SuperM2M on real data, as it decouples enhancement and ASR and does not employ joint training.

### v3

@@ 7.7.1 SNR Augmentation for Simulated Training Mixtures
7.7.2 iSTFT-STFT Projection
We apply inverse STFT (iSTFT) followed by STFT operations to the DNN estimates $\hat{X}_{q}$ and $\hat{V}_{q}$ before loss computation, i.e., $\hat{X}_{q}\coloneqq\text{STFT}(\text{iSTFT}(\hat{X}_{q}))$ and $\hat{V}_{q}\coloneqq\text{STFT}(\text{iSTFT}(\hat{V}_{q}))$. See Fig. 1 for an illustration. This often yields slight improvement, as the losses now penalize the RI components and magnitudes extracted from re-synthesized signals, which are the final system output used for human hearing and downstream tasks [82, 83, 69]. Notice that, in Fig. 2, ASR features are extracted from re-synthesized signals.
7.7.3 Run-Time Speaker Reinforcement for Robust ASR
At run time, in default we feed $\hat{x}_{q}$ for ASR. Alternatively, we employ a technique named speaker reinforcement [72], where $\hat{x}_{q}$ is re-mixed with the input mixture $y_{q}$ at an energy level of $\gamma$ dB before recognition. See Fig. 2 for an illustration. That is, $\hat{x}_{q}+\eta\times y_{q}$, where $\eta\in{\mathbb{R}}_{>0}$ and $\gamma=10\times\text{log}_{10}(\|\hat{x}_{q}\|^{2}/\|\eta\times y_{q}\|^{2})$. We find this technique usually effective for ASR, as the re-mixed input mixture can alleviate distortion to target speech.
8 Evaluation Results
This section reports our evaluations results on CHiME-4. Following earlier studies [35, 40], we set $I_{p}$ to $20$ for all the non-reference far-field microphones and the close-talk microphone, and $J_{p}$ to $1$ for all the non-reference far-field microphones. $J_{0}$, the future filter tap for the close-talk microphone, can be tuned to a value shared by all the training mixtures, or it can be estimated, on the fly, for each training mixture by using the method described in Section 5. All the six far-field microphones are used for computing $\mathcal{L}_{\text{MC}}$ in (4). We emphasize that we spent minimal amount of effort on hyper-parameter tuning. Other hyper-parameter setup could lead to better performance.
Table 3: SuperM2M vs. purely-supervised models on CHiME-4 (#input mics: $1$; ASR model: Whisper Large v2)
[…]
2 Supervised S iNeuBe [13] - - - - - $15.1$ $0.954$
4e SuperM2M S+R TFGridNetv1 5 6+1 - 16.901 048 553 260 8916.9010485532608916.901\,048\,553\,260\,8916.901 048 553 260 89 17.357 574 660 580 04517.35757466058004517.357\,574\,660\,580\,04517.357 574 660 580 045 2.362 516 766 335 0252.3625167663350252.362\,516\,766\,335\,0252.362 516 766 335 025 0.959 485 302 685 890 20.95948530268589020.959\,485\,302\,685\,890\,20.959 485 302 685 890 2 7.341 077 767 021 627.341077767021627.341\,077\,767\,021\,627.341 077 767 021 62 5.288 209 430 841 8385.2882094308418385.288\,209\,430\,841\,8385.288 209 430 841 838 12.226 013 234 077 7512.2260132340777512.226\,013\,234\,077\,7512.226 013 234 077 75 7.507 367 767 954 0877.5073677679540877.507\,367\,767\,954\,0877.507 367 767 954 087
4f SuperM2M S+R TFGridNetv1 6 6+1 - 16.859 340 666 821 5116.8593406668215116.859\,340\,666\,821\,5116.859 340 666 821 51 17.379 205 978 436 4717.3792059784364717.379\,205\,978\,436\,4717.379 205 978 436 47 2.414 184 342 550 5092.4141843425505092.414\,184\,342\,550\,5092.414 184 342 550 509 0.961 652 581 435 139 40.96165258143513940.961\,652\,581\,435\,139\,40.961 652 581 435 139 4 7.341 077 767 021 627.341077767021627.341\,077\,767\,021\,627.341 077 767 021 62 5.288 209 430 841 8385.2882094308418385.288\,209\,430\,841\,8385.288 209 430 841 838 12.226 013 234 077 7512.2260132340777512.226\,013\,234\,077\,7512.226 013 234 077 75 7.507 367 767 954 0877.5073677679540877.507\,367\,767\,954\,0877.507 367 767 954 087
4g SuperM2M S+R TFGridNetv1 7 6+1 - 16.963 124 431 443 93716.96312443144393716.963\,124\,431\,443\,93716.963 124 431 443 937 17.398 348 091 374 1717.3983480913741717.398\,348\,091\,374\,1717.398 348 091 374 17 2.420 927 345 301 165 72.42092734530116572.420\,927\,345\,301\,165\,72.420 927 345 301 165 7 0.961 293 461 207 607 60.96129346120760760.961\,293\,461\,207\,607\,60.961 293 461 207 607 6 7.228 138 109 067 4417.2281381090674417.228\,138\,109\,067\,4417.228 138 109 067 441 5.021 983 784 437 9015.0219837844379015.021\,983\,784\,437\,9015.021 983 784 437 901 12.194 995 864 350 70312.19499586435070312.194\,995\,864\,350\,70312.194 995 864 350 703 6.902 435 241 197 4566.9024352411974566.902\,435\,241\,197\,4566.902 435 241 197 456
4h SuperM2M S+R TFGridNetv1 8 6+1 - 16.627 877 601 619 9216.6278776016199216.627\,877\,601\,619\,9216.627 877 601 619 92 17.397 258 721 839 1217.3972587218391217.397\,258\,721\,839\,1217.397 258 721 839 12 2.414 387 100 934 9822.4143871009349822.414\,387\,100\,934\,9822.414 387 100 934 982 0.960 830 590 588 033 90.96083059058803390.960\,830\,590\,588\,033\,90.960 830 590 588 033 9 7.260 406 582 768 6357.2604065827686357.260\,406\,582\,768\,6357.260 406 582 768 635 5.026 017 506 353 1125.0260175063531125.026\,017\,506\,353\,1125.026 017 506 353 112 12.045 078 577 336 64112.04507857733664112.045\,078\,577\,336\,64112.045 078 577 336 641 6.954 138 875 962 986.954138875962986.954\,138\,875\,962\,986.954 138 875 962 98
[…]
7a SuperM2M S+R TFGridNetv2 est. 6+1 - 17.287 956 691 601 0617.2879566916010617.287\,956\,691\,601\,0617.287 956 691 601 06 17.666 113 485 961 7117.6661134859617117.666\,113\,485\,961\,7117.666 113 485 961 71 2.443 089 644 385 106 72.44308964438510672.443\,089\,644\,385\,106\,72.443 089 644 385 106 7 0.963 246 019 069 4850.9632460190694850.963\,246\,019\,069\,4850.963 246 019 069 485 6.776 379 477 250 725 56.77637947725072556.776\,379\,477\,250\,725\,56.776 379 477 250 725 5 4.933 241 902 303 2554.9332419023032554.933\,241\,902\,303\,2554.933 241 902 303 255 11.424 731 182 795 69811.42473118279569811.424\,731\,182\,795\,69811.424 731 182 795 698 6.307 843 441 393 936.307843441393936.307\,843\,441\,393\,936.307 843 441 393 93
7b SuperM2M S+R TFGridNetv2 est. 6+1 - 17.101 209 408 406 05617.10120940840605617.101\,209\,408\,406\,05617.101 209 408 406 056 17.505 491 557 785 8117.5054915577858117.505\,491\,557\,785\,8117.505 491 557 785 81 2.428 467 516 014 072.428467516014072.428\,467\,516\,014\,072.428 467 516 014 07 0.961 963 570 948 513 70.96196357094851370.961\,963\,570\,948\,513\,70.961 963 570 948 513 7 7.022 426 589 222 3317.0224265892223317.022\,426\,589\,222\,3317.022 426 589 222 331 4.739 623 250 373 119 54.73962325037311954.739\,623\,250\,373\,119\,54.739 623 250 373 119 5 11.243 796 526 054 5911.2437965260545911.243\,796\,526\,054\,5911.243 796 526 054 59 6.230 287 989 245 6446.2302879892456446.230\,287\,989\,245\,6446.230 287 989 245 644
7c SuperM2M S+R TFGridNetv2 est. 6+1 - 16.605 623 853 251 7916.6056238532517916.605\,623\,853\,251\,7916.605 623 853 251 79 17.418 666 160 478 68317.41866616047868317.418\,666\,160\,478\,68317.418 666 160 478 683 2.510 620 386 221 1922.5106203862211922.510\,620\,386\,221\,1922.510 620 386 221 192 0.962 660 024 574 209 50.96266002457420950.962\,660\,024\,574\,209\,50.962 660 024 574 209 5 7.046 627 944 498 2257.0466279444982257.046\,627\,944\,498\,2257.046 627 944 498 225 4.783 994 191 440 4424.7839941914404424.783\,994\,191\,440\,4424.783 994 191 440 442 11.342 018 196 856 90711.34201819685690711.342\,018\,196\,856\,90711.342 018 196 856 907 6.023 473 450 183 5486.0234734501835486.023\,473\,450\,183\,5486.023 473 450 183 548
[…]
9
Close-talk
Mixture
@@ 8 Evaluation Results
Table 7.7.1 and 8 respectively report $6$- and $1$-channel enhancement performance on the fifth microphone of the CHiME-4 simulated test data, and robust ASR performance on the official CHiME-4 test utterances (by using the Whisper Large v2 model for recognition). The two tables consist of exactly the same set of experiments, and differ only in the number of input microphones to the enhancement models.
On the simulated test data, purely supervised models (in system $1$a-$1$b) trained on the official simulated training data (denoted as S in the “Training data” column) produce large improvement over unprocessed mixtures (e.g., in system 1a of Table 7.7.1, $22.2$ vs. $7.5$ dB SISDR, and in 1a of Table 8, $17.1$ vs. $7.5$ dB SISDR), and achieve enhancement performance better than existing supervised models such as iNeuBe (based on TCN-DenseUNet) [13, 84], SpatialNet [23], USES [24] and USES2 [25] in both $1$- and $6$-channel cases. However, the ASR performance is much worse than directly using unprocessed mixtures for ASR, especially in multi-channel cases (e.g., in system $1$a and 00 of Table 7.7.1, $53.23$% vs. $7.69$% WER on REAL test data, and in system $1$a and 00 of Table 8, $10.20$% vs. $7.69$% WER on REAL test data). This degradation has been widely observed by existing studies [33, 43], largely due to the mismatches between simulated training and real-recorded test conditions, and the speech distortion incurred by enhancement.
In system $7$c of Table 7.7.1 and $7$b of Table 8, we respectively show the results of our best performing SuperM2M models (without employing the speaker reinforcement technique described in Section 7.7.3) for $6$- and $1$-channel cases. By training on the official simulated and real data combined (denoted as S+R), SuperM2M obtains clearly better ASR results on the real test data than the purely-supervised models (in system $1$a-$1$b) and unprocessed mixtures (in system 00), and the enhancement performance on the simulated test data remains strong. These results indicate that SuperM2M can effectively learn from real data, has better generalizability to real data, and can perform enhancement with low distortion to target speech and high reduction to non-target signals.
@@ 8.2 Effects of Including Loss on Close-Talk Mixture in $\mathcal{L}_{\text{MC}}$
First, in system $3$, we report the results of not including close-talk microphone in the loss function. That is, in (4) we do not include the $\mathcal{L}_{\text{MC},0}$ loss, when training SuperM2M. We observe clear improvement in ASR performance on the real test data over the purely-supervised model in system $1$a (i.e., $4.46$% vs. $53.23$% in Table 7.7.1 and $7.02$% vs. $10.20$% in Table 8), indicating that close-talk mixtures are not must-have for our system.
Next, we include close-talk mixtures for training. In system $4$a-$4$h, the number of future taps for the close-talk microphone, $J_{0}$, is tuned to a value shared by all the training utterances, and we enumerate $J_{0}$ from $1$ all the way up to $8$. In system $5$, we instead estimate $J_{0}$, on the fly, for each training mixture during training, using the technique described in Section 5. Compared with system $4$a-$4$h, system $5$ obtains better ASR performance on the real test data, indicating the effectiveness of estimating $J_{0}$.
In both Table 7.7.1 and 8, comparing system $5$ and $3$, we observe that including close-talk mixtures for training SuperM2M produces better ASR performance on the real test data of CHiME-4 (e.g., $3.97$% vs. $4.46$% in Table 7.7.1 and $6.51$% vs. $7.02$% in Table 8).
@@
In system $6$a-$6$d of Table 7.7.1 and 8, we respectively report the results of UNSSOR [35] and M2M [40]. UNSSOR is trained in an unsupervised way directly on far-field mixtures by using the $\mathcal{L}_{\text{MC}}$ loss in (4) but without including $\mathcal{L}_{\text{MC},0}$ (i.e., the loss on close-talk mixture). M2M is the same as SuperM2M but without the supervised learning branch. We can train UNSSOR and M2M on the real data alone (denoted as R), or on the simulated and real data combined (i.e., S+R) considering that SuperM2M is trained on S+R. When M2M is trained on S+R, the loss function on the simulated data does not include the loss on close-talk mixture, since CHiME-4 does not provide simulated close-talk mixtures.
@@
Comparing system $6$a-$6$d and $5$, we can see that SuperM2M obtains much better performance than UNSSOR and M2M, suggesting the effectiveness of the proposed co-learning mechanism. The low performance of M2M and UNSSOR is possibly because the noise in CHiME-4 typically consists of an unknown number of diffuse and directional sources. In this case, unsupervised algorithms with two hypothesized directional sources tend to get confused about which source should be the target speech source. This problem is naturally avoided by the supervised learning mechanism in SuperM2M, which models noise sources as a single, combined source.
@@ 8.4 Miscellaneous Results
In system $7$a, $7$b, and $7$c of Table 7.7.1 and 8, we switch TF-GridNet from v1 to v2, perform iSTFT-STFT projection, and apply SNR augmentation. Comparing system $7$a with $7$b, we observe that applying iSTFT-STFT projection can improve ASR performance in most cases. From system $5$ to $7$c, although the ASR performance on the real test data gets better in Table 8 while worse in Table 7.7.1, in both tables better ASR performance is observed on the real validation set. We therefore use system $7$c as our default system for the rest experiments of this paper.
@@
System Approach Frontend SIMU REAL SIMU REAL
@@
2b 7c of Table 8 SuperM2M TFGridNetv2 1 $10$ 2.400 442 4772.4004424772.400\,442\,4772.400 442 477 1.640 915 9621.6409159621.640\,915\,9621.640 915 962 4.538 662 6824.5386626824.538\,662\,6824.538 662 682 2.400 859 4512.4008594512.400\,859\,4512.400 859 451
@@
4b 7c of Table 7.7.1 SuperM2M TFGridNetv2 2 $10$ 1.283 185 841.283185841.283\,185\,841.283 185 84 1.331 170 0281.3311700281.331\,170\,0281.331 170 028 1.881 770 6381.8817706381.881\,770\,6381.881 770 638 1.835 676 5841.8356765841.835\,676\,5841.835 676 584
@@
6b 7c of Table 7.7.1 SuperM2M TFGridNetv2 6 $10$ 0.833 333 3330.8333333330.833\,333\,3330.833 333 333 1.231 608 8351.2316088351.231\,608\,8351.231 608 835 1.372 805 3791.3728053791.372\,805\,3791.372 805 379 1.578 775 2811.5787752811.578\,775\,2811.578 775 281
@@
Notes: The best scores are highlighted in bold in the $1$-, $2$- and $6$-channel cases separately.
@@ 8.6 Breaking Out to New Highs on CHiME-4 ASR Tasks
Table 8.5 reports the ASR performance of SuperM2M based on the official ASR setup of CHiME-4, using the ASR model proposed in [53] and still following the evaluation pipeline in Fig. 2. Comparing system 00b with 00a, we observe that we have successfully reproduced the ASR system proposed in [53].
SuperM2M, despite not jointly trained with the ASR model, achieves a new state-of-the-art on the REAL test set in each of the $1$-, $2$- and $6$-channel tasks, significantly outperforming the previous best obtained by IRIS [53] and MultiIRIS [87] (e.g., in the $1$-channel case $3.04$% vs. $3.92$% WER in $2$a and $1$b, in the $2$-channel case $1.94$% vs. $2.65$% WER in $4$a and $3$b, and in the $6$-channel case $1.61$% vs. $1.77$% WER in $6$a and $5$b). IRIS jointly trains a Conv-TasNet based monaural enhancement model, a WavLM based ASR feature extractor, and an encoder-decoder transformer based ASR model. MultiIRIS, building upon IRIS, replaces the Conv-TasNet module with a DNN based weighted power minimization distortionless response (WPD) beamformer. From system $1$a vs. $1$b, $3$a vs. $3$b, and $5$a vs. $5$b, we observe that, without joint training, IRIS and MultiIRIS often obtain clearly worse ASR performance, especially in the $1$- and $2$-channel cases. These results further indicate the effectiveness and potential of SuperM2M on real data, as it decouples enhancement and ASR and does not employ joint training.